A sound wave traveling at 342 m/s is emitted
by the foghorn of a tugboat. An echo is heard
3.60 s later.
How far away is the reflecting object?
Answer in units of m.
1
Expert's answer
2011-02-03T05:11:38-0500
The path of the wave is
S = Vt = 342*3.6 = 1231.2 m
The distance from the reflecting object would be
D = S/2 = 615.6 m